# Language Models as Causal Effect Generators

### v1

@@ 4.1 Breast cancer SD-SCM
4.2 Effect estimation results
We compare the performance of several treatment effect estimation algorithms on this example benchmark. As a naive baseline, ordinary least squares using only the treatment $\tilde{t}$ is considered (T-Only OLS). Against this baseline, we consider several causal inference methods of different types, including the causal forest [56, 3] (CausalForest) and two double machine learning methods for CATE estimation, one linear (LinearDML) and one non-parametric (ForestDML) [9, 3, 46, 8, 19, 43, 5]. For comparison we include also two doubly robust meta-learning methods [35], again, one linear (LinearDR) and one non-parametric (ForestDR). We add also Bayesian additive regression trees (BART) [27, 11] as a widely-used Bayesian non-parametric example. As points of reference for NN-based CATE estimation methods, we include an NN-based T-learner (TNet), and the NN-based TARNet [53]. Additional baselines include a random forest baseline (RF) that fits a single response surface and directly predicts treatment effects for each unit, and a linear regression baseline (LinReg) that takes the conditional mean difference (the fit coefficient on $\tilde{t}$) to be the effect. Finally, we include two methods that target ITEs specifically. One method uses BART posterior draws specifically for ITEs instead of CATEs (BART-ITE), and the other is conformalized counterfactual quantile regression (CQR) [39], which provides conformal inference-based interval estimates of ITEs.
All methods are fit using the default settings of their publicly-available implementations. The causal forest, DML, and DR implementations are provided by [5], the BART methods by [16], the NN-based methods by [12, 14, 13] and CQR by [39]. [^7] All code to reproduce this benchmark is available at https://github.com/lbynum/sequence-driven-scms. While additional hyperparameter tuning and model optimization could be performed for several methods on a case-by-case basis, this section demonstrates what estimation results we get off-the-shelf.
4.2.1 Average treatment effects
Method $\mathcal{P}=$ GPT-2 $\mathcal{P}=$ Llama-3-8b
$R^{2}$ RMSE $R^{2}$ RMSE
All Cov. Hidden All Cov. Hidden All Cov. Hidden All Cov. Hidden
T-Only OLS 0.6047 0.6047 0.0172 0.0172 0.5082 0.5082 0.0091 0.0091
BART 0.9999 0.8794 0.0003 0.0095 0.9967 0.8476 0.0007 0.0051
ForestDML 0.9941 0.8686 0.0021 0.0099 0.9608 0.8129 0.0026 0.0056
ForestDR $\leq 0$ $\leq 0$ 6.4268 29.0210 0.9581 0.8179 0.0027 0.0055
LinReg 0.9146 0.8646 0.0080 0.0101 0.6538 0.5599 0.0076 0.0086
LinearDML 0.979 0.8655 0.0040 0.0100 0.9608 0.8216 0.0026 0.0055
LinearDR $\leq 0$ $\leq 0$ 11.0736 29.5414 0.9589 0.8176 0.0026 0.0055
RF 0.976 0.8766 0.0042 0.0096 0.8122 0.8295 0.0056 0.0054
For all implementations that directly support ATE estimation, we report the $R^{2}$ and root-mean-squared-error (RMSE) across the 1000 datasets for each language model in two settings, using $\text{logP}(\tilde{y}=0)$ as the outcome. The first setting is with estimation using all 14 covariates (all 12 confounders, the treatment, and the outcome). This is denoted All Cov. in Table 1. The second setting is with the variables $\mathbf{U}=\{\tilde{u}_{1},\tilde{u}_{2},\tilde{u}_{3},\tilde{u}_{4}\}=$age, medical conditions, medication, menopausal status} hidden, denoted Hidden. Table 1 shows that ATE estimation is more or less challenging depending on which language model is used. In this case, Llama-3-8b produces ATEs that are more challenging to estimate, with the exception of GPT-2 for the doubly robust methods, whose $R^{2}$ and RMSE suffer significantly due to several large outlying estimates. Across all methods, performance tends to drop significantly in the ‘Hidden’ setting, suggesting that $\mathbf{U}$ are indeed hidden confounders. We refer to this setting henceforth as ‘Hidden Confounding.’ Across methods, BART shows the strongest performance in all settings in Table 1.
As discussed in Section 1, a particular benefit of SD-SCM-based data generation is access to individual-level potential outcomes rather than ATEs only, allowing for comparisons of performance on CATE and ITE estimation.
4.2.2 Conditional and individual treatment effects
Figure 5: $R^{2}$ values (clipped at zero) across all methods that estimate CATEs.
Figure 5 shows $R^{2}$ values (clipped at zero) across all methods that provide point estimates for CATEs. When all covariates are observed, BART does significantly better explaining CATE variation, while DML and DR methods do as well at times but with a much lower average. However, CATE estimation becomes much more challenging for all methods under hidden confounding. The difference in effect estimation difficulty between Llama-3-8b and GPT-2 is also more noticeable for CATEs than it was for ATEs. Figures 9 and 10 in Appendix B show the same results in terms of PEHE (Precision in Estimating Heterogeneous Effects) [27], revealing that with no clipping, DR methods show large outliers with GPT-2, and NN methods show large outliers instead with Llama-3-8b.
Figure 6: CATE empirical coverage ($\alpha=0.05$) and interval width (in outcome standard deviation units) for methods that provide intervals. Nominal coverage of 95% is indicated by the red line.
Figure 7: Interval estimates from BART versus BART-ITE across two example datasets, (a) and (b).
When the ITE varies due to any covariates not conditioned on in the CATE, the two quantities are distinct. Thus, the Hidden Confounding setting for CATE estimation shows us how CATE estimators perform when the target is instead the ITE. In this setting, where estimation is difficult and no methods perform well, uncertainty is especially important. Figure 6 shows empirical coverage results for all estimators that provide intervals. Empirical coverage is under nominal for all methods that target CATE in the setting with all covariates. Hidden confounding generally increases uncertainty, but this does make coverage closer to nominal, especially for LinearDML and LinearDR. CQR remains at or above nominal coverage, but with much wider intervals than the other methods. Interestingly, BART for CATE achieves higher median coverage of the ITE than BART-ITE, but with a much larger tail of poor coverage. BART-ITE, by contrast, has much less variable coverage in the ITE setting, but at the cost of much wider intervals. Figure 7 demonstrates this further with intervals for BART targeting the CATE versus BART-ITE across two example datasets. When all covariates are observed (top row), there are cases where intervals from either method are informative about individual-level effects. However, under hidden confounding (bottom row), the tighter intervals of BART targeting the CATE are overconfident with variable coverage, and the wider intervals of BART-ITE are so wide as to be vacuous, even if we want to recover just the ranking of the ITEs.
Takeaways.
We summarize a few takeaways for off-the-shelf estimation performance with these example SD-SCMs. The first is that less complicated methods are often able to perform well. Second, in a real-world setting, where a method might often be used with its default parameters, stability can be important (e.g., the DR and NN method performance suffers often due to lack of stability). The third takeaway is that hidden confounding has a big impact, across all methods. Even methods that perform particularly well with all covariates (like BART) suffer significantly under hidden confounding. Finally, ITE intervals can be unstable and/or vacuous for decision making, especially under hidden confounding, and should thus be used carefully.
5 Auditing language models for (un)desirable causal effects
The same framework we use to generate causal effects and benchmark effect estimation methods can allow us to inspect what causal information has been encoded semantically in an LLM. Our collection of breast cancer SD-SCMs are already set up to explore the effect of PD-L1 on chosen therapy plans, while allowing us to marginalize out an important source of variability: phrasing. Essentially, this amounts to reverse engineering the decision-making process of clinicians, as learned
from whatever data the language model were trained on. Recall from Section 4.1 that there are nine possible outcomes we could explore in the breast cancer example. Figure 8 shows one example where the two language models strongly disagree on what the causal effect is. The effect in this case is the change in probability of choosing the therapy plan “start a regimen of trastuzumab and pertuzumab” (shown in standard deviation units). GPT-2 has encoded that on average, an increase in PD-L1 expression levels has neither a positive nor negative impact on whether or not this therapy will be chosen. However, Llama-3-8b has encoded instead that an increase in PD-L1 always increases the likelihood of this therapy plan. This discrepancy indicates that these two language models have encoded two meaningfully different causal models. We believe this same procedure can underpin more thorough auditing of LLMs for misinformation or discrimination, enabling, for example, path-specific counterfactual fairness analysis [36, 10].
6 Conclusion and Future Work
In this work, we have introduced sequence-driven structural causal models (SD-SCMs)as a framework for LLM-based data generation with controllable causal structure, allowing for observational, interventional, and counterfactual data generation according to user-defined DAGs and LLM-defined structural equations. We demonstrate an important use-case for SD-SCMs by creating a benchmark for causal effect estimation. In this proof of concept, we focused on methods that assume observed confounders, but there are many other settings to explore for effect estimation, such as instrumental variables [1, 26]. Using SD-SCMs to additionally test causal discovery is of immediate interest, for example, allowing us to test whether a structure learning method can identify whether one variable is causally upstream or downstream of another [34]. Another significant area of future work is to use SD-SCMs or similar as a means of specifying causal structure over sequential data during learning [29]; rather than use an already-trained language model to generate effects, a model can be trained or fine-tuned to handle tasks that require causal reasoning, including complex confounding and sequential decision making.
In short, we believe SD-SCMs can serve as a stepping stone for any application that would benefit from sequential data with controllable causal structure, opening the door to several types of applications.
Acknowledgments
This research was supported in part by National Science Foundation (NSF) award No. 1922658 and the Samsung Advanced Institute of Technology (under the project Next Generation Deep Learning: From Pattern Recognition to AI). We are grateful to Roman Mutel and Yulia Maksymiuk for involvement in early discussions of this work and especially grateful to Julia Stoyanovich for helpful feedback on data generation.
Appendix A Full description of the breast cancer SD-SCMs
@@
$(\tilde{u}_{1},\ldots,\tilde{u}_{4},\tilde{x}_{1},\ldots,\tilde{x}_{8},\tilde{t},\tilde{y})=(4,2,0,1,3,2,3,0,2,0,1,1,0,3).$
@@
The full set of possible examples and code to generate this SD-SCM is available in our repository at https://github.com/lbynum/sequence-driven-scms.
Appendix B Additional estimation results
@@
Figure 9: PEHE across all methods that estimate CATE, shown in standard deviation units of the outcome $\tilde{y}$.
Figure 10: PEHE across all methods that estimate CATE, shown in units of ITE standard deviation.
Figure 9 shows standardized Precision in Estimating Heterogeneous Effects (PEHE) [27], which is the RMSE of the CATE predictions across the different observed values of $x$, i.e.,
@@
[…]
Figure 10 shows the same metric standardized instead using the (much smaller) standard deviation of the ITE.

### v2

@@ Appendix B Full description of the breast cancer SD-SCMs
[…]
$(4,2,0,1,3,2,3,0,2,0,1,1,0,3).$
@@
The full set of possible examples and code to generate this SD-SCM and corresponding data (in our case generated using V100 and RTX8000 GPUs) is available in our repository at https://github.com/lbynum/sequence-driven-scms.
Method $\mathcal{P}=$ GPT-2 $\mathcal{P}=$ Llama-3-8b
$R^{2}$ RMSE $R^{2}$ RMSE
All Cov. Hidden All Cov. Hidden All Cov. Hidden All Cov. Hidden
T-Only OLS 0.6047 0.6047 0.0172 0.0172 0.5082 0.5082 0.0091 0.0091
BART 0.9999 0.8794 0.0003 0.0095 0.9967 0.8476 0.0007 0.0051
ForestDML 0.9941 0.8686 0.0021 0.0099 0.9608 0.8129 0.0026 0.0056
ForestDR $\leq 0$ $\leq 0$ 6.4268 29.0210 0.9581 0.8179 0.0027 0.0055
ForestS 0.9771 0.8777 0.0041 0.0096 0.8243 0.8286 0.0054 0.0054
ForestT 0.9793 0.8588 0.0039 0.0103 0.9454 0.8126 0.0030 0.0056
LinReg 0.9146 0.8646 0.0080 0.0101 0.6538 0.5599 0.0076 0.0086
LinearDML 0.979 0.8655 0.0040 0.0100 0.9608 0.8216 0.0026 0.0055
LinearDR $\leq 0$ $\leq 0$ 11.0736 29.5414 0.9589 0.8176 0.0026 0.0055
LinearS 0.9146 0.8632 0.0080 0.0101 0.6538 0.4869 0.0076 0.0093
LinearT 0.9181 0.8688 0.0078 0.0099 0.6395 0.5385 0.0078 0.0088
RF 0.976 0.8766 0.0042 0.0096 0.8122 0.8295 0.0056 0.0054
Table 2: Mean $R^{2}$ of methods estimating CATEs/ITEs, comparing estimation with datasets of size 1,000 versus 10,000.
Method All Covariates Hidden $\mathbf{U}$
$\mathcal{P}=$ GPT-2 $\mathcal{P}=$ Llama-3-8b $\mathcal{P}=$ GPT-2 $\mathcal{P}=$ Llama-3-8b
1,000 10,000 1,000 10,000 1,000 10,000 1,000 10,000
T-Only OLS $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0
BART 0.7162 0.9691 0.5221 0.9214 $\leq$0 $\leq$0 $\leq$0 $\leq$0
BART-ITE 0.7102 0.9344 0.5183 0.8823 0.0054 0.0185 0.0011 0.0169
CausalForest $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0 0.0313 $\leq$0 0.1399
ForestDML $\leq$0 0.6605 $\leq$0 0.5551 $\leq$0 0.0608 $\leq$0 0.1850
ForestDR $\leq$0 0.7220 $\leq$0 0.5968 $\leq$0 0.0503 $\leq$0 0.1931
ForestS $\leq$0 0.0661 $\leq$0 0.1263 0.0026 0.0371 0.0063 0.0936
ForestT 0.1031 0.3202 $\leq$0 0.1319 0.0076 0.0366 0.0054 0.0719
LinReg $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0
LinearDML 0.2381 0.5309 0.1491 0.5404 $\leq$0 $\leq$0 $\leq$0 0.1483
LinearDR 0.2329 0.5900 0.1562 0.5293 $\leq$0 $\leq$0 $\leq$0 0.1734
LinearS $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0
LinearT 0.0116 0.1470 $\leq$0 0.1355 $\leq$0 $\leq$0 $\leq$0 $\leq$0
RF $\leq$0 0.0327 $\leq$0 0.1038 0.0024 0.0373 0.0060 0.0940
TARNet $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0
TNet $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0 $\leq$0
Figure 8: Change in mean $R^{2}$ (if above 0) of methods estimating CATEs and ITEs after a tenfold increase in dataset size from $1,000$ to $10,000$.
Figure 9: $R^{2}$ values across all methods that provide point estimates of CATEs/ITEs for datasets of size 10,000 generated by GPT-2 (left) and Llama-3-8b (right).
Figure 10: CATE/ITE empirical coverage ($\alpha=0.05$) and interval width (in outcome standard deviation units) for methods that provide intervals. Nominal coverage of 95% is indicated by the red line. Shown for datasets of size 10,000 generated by GPT-2 (left) and Llama-3-8b (right).
@@
Figure 11: PEHE across all methods that provide point estimates of CATEs/ITEs, shown in standard deviation units of the outcome $\tilde{y}$. Results shown for datasets of size 10,000 generated by GPT-2 (left) and Llama-3-8b (right).
Figure 12: PEHE across all methods that provide point estimates of CATEs/ITEs, shown in units of ITE standard deviation. Results shown for datasets of size 10,000 generated by GPT-2 (left) and Llama-3-8b (right).
Figure 13: $R^{2}$ values across all methods that provide point estimates of CATEs/ITEs for datasets of size 1,000 generated by GPT-2 (left) and Llama-3-8b (right).
Figure 14: CATE/ITE empirical coverage ($\alpha=0.05$) and interval width (in outcome standard deviation units) for methods that provide intervals. Nominal coverage of 95% is indicated by the red line. Shown for datasets of size 1,000 generated by GPT-2 (left) and Llama-3-8b (right).
Figure 15: Interval estimates from BART versus BART-ITE across two example datasets of size 1,000, (a) and (b).
Figure˜7 shows plots of the features (LABEL:fig:bcancer_hist), propensity scores (LABEL:fig:bcancer_propensity), categorical ITEs (LABEL:fig:bcancer_ites_binary), and continuous ITEs (LABEL:fig:bcancer_ites) for a single generated dataset using Llama-3-8b. Because the outcome $\tilde{y}$ has $|\Omega_{\tilde{y}}|=4$ possible values, we can consider several possible outcomes, including the observed outcome (categorical), log probabilities for each outcome value (continuous), or probabilities for each outcome value (continuous). This creates, in effect, nine possible targets for each dataset. For benchmarking purposes, we find using probabilities and/or log probabilities as the outcome to be the most useful — there is frequently an effect (even at the individual level) in probability space, even if the sampled outcomes do not change. Comparing Figure LABEL:fig:bcancer_ites_binary to Figure LABEL:fig:bcancer_ites demonstrates this, where for the 400 or so observations where the categorical ITE is zero, the continuous ITE is instead nonzero.
Figure LABEL:fig:bcancer_ites also demonstrates that we are able to satisfy our main criterion for meaningful benchmark: the observational distribution $P_{\tilde{y}\mid\tilde{t}=t}^{\mathfrak{B}}$ and interventional distribution $P_{\tilde{y}}^{\mathfrak{B};\text{do}(\tilde{t}=t)}$ are different enough that the SATE and the observed mean difference in outcomes between the treatment and control group are not only different in value, but they also disagree in sign. This is particularly meaningful in a causal inference setting — the treatment appears to lower the outcome, when in fact, its effect is to increase the outcome.
Appendix C Additional estimation results
C.1 ATE results
For all implementations that directly support ATE estimation, we report the $R^{2}$ and root-mean-squared-error (RMSE) across the 1000 datasets for each language model in two settings, using $\text{logP}(\tilde{y}=0)$ as the outcome. The first setting is with estimation using all 14 covariates (all 12 confounders, the treatment, and the outcome). This is denoted All Cov. in Table˜1. The second setting is with the variables $\mathbf{U}=\{\tilde{u}_{1},\tilde{u}_{2},\tilde{u}_{3},\tilde{u}_{4}\}=$ {age, medical conditions, medication, menopausal status} hidden, denoted Hidden. Table˜1 shows that ATE estimation is more or less challenging depending on which language model is used. In this case, Llama-3-8b produces ATEs that are more challenging to estimate, with the exception of GPT-2 for the doubly robust methods, whose $R^{2}$ and RMSE suffer significantly due to several large outlying estimates. Across all methods, performance tends to drop significantly in the ‘Hidden’ setting, suggesting that $\mathbf{U}$ are indeed hidden confounders. Across methods, BART shows the strongest performance in all settings in Table˜1.
C.2 Comparison of dataset size 1,000 and 10,000
Tables˜2 and 8 show that CATE and ATE estimation remain difficult even after a tenfold increase in dataset size (from $N=1,000$ to $N=10,000$), especially in the Hidden $\mathbf{U}$ setting. Across estimation methods, performance tends to increase as sample size increases, especially if the method originally achieved $R^{2}$ above zero with $N=1,000$. In other words, methods that do reasonably well at $N=1,000$ show improvement with more data, as we would expect. However, several methods struggle in both settings, even with ten times more data. For example, TARNet, TNet, and CausalForest still remain unstable and inaccurate in both the All Covariates and the Hidden $\mathbf{U}$ settings across both sample sizes. Overall, these results indicate that CATE and ITE estimation in this case are not challenging due only to small sample sizes. This is useful to know, especially when we consider that corresponding real-world use-cases often deal with even smaller sample sizes.
Additional results for each dataset size are shown individually in the following figures. Figures˜9 and 10 show $R^{2}$ and coverage results on datasets of size 10,000. These correspond to the same figures in the main text, but now showing both GPT-2 and Llama-3-8b, allowing for comparison across models. Figure˜11 shows the same setting using standardized Precision in Estimating Heterogeneous Effects (PEHE) Hill (2011), which is the RMSE of the CATE predictions across the different observed values of $x$, i.e.,
@@ C.2 Comparison of dataset size 1,000 and 10,000
$\text{(Standardized PEHE)}_{j}=\sqrt{\frac{1}{n\cdot\hat{\sigma}^{2}_{j}}\sum_{i=1}^{n}(\hat{\tau}_{i}-\tau_{i})^{2}}.$
Figure˜12 shows the same metric standardized instead using the (much smaller) standard deviation of the ITE.
Results in the case of dataset size $1,000$ show similar trends to those in the size $10,000$ setting. Figure˜13 shows $R^{2}$ values clipped at zero across all methods that provide point estimates for CATEs. When all covariates are observed, BART does significantly better explaining CATE variation, followed by DML and DR with much lower averages, much like the size 10,000 case. Similarly, CATE estimation becomes much more challenging for all methods with hidden $\mathbf{U}$. The difference in effect estimation difficulty between Llama-3-8b and GPT-2 is also more noticeable for CATEs than it was for ATEs. Overall, some methods show more instability in the dataset size 1,000 case than in the size 10,000 case, as expected with less data.
Figure˜14 shows empirical coverage results in the dataset size 1,000 case for all estimators that provide intervals. Similar to the size 10,000 case, empirical coverage is under nominal for all methods that target CATE in the setting with all covariates. Hidden $\mathbf{U}$ generally increases uncertainty, bringing the DR methods and LinearDML median coverage near nominal. Interestingly, BART for CATE achieves higher median coverage of the ITE than BART-ITE, but with a much larger tail of poor coverage. BART-ITE, by contrast, has much less variable coverage in the ITE setting, but at the cost of much wider intervals. Figure˜15 shows intervals for BART targeting the CATE versus BART-ITE across two example datasets of size 1,000, demonstrating that, as in the size 10,000 case, the tighter intervals of BART targeting the CATE can be overconfident with variable coverage, while the wider intervals of BART-ITE are too wide to be useful.